# Galled Perfect Transfer Networks

### v2

@@ 3 The Galled Completion Problem
Let $T$ be a tree on leafset $\mathcal{S}$ and character set $\mathcal{C}$. If $T$ is galled-completable, then in the redundancy-free network $N=(V_{T},E_{T})$ of $T$, every node is incident to at most one transfer edge.
@@
Suppose that some node $u^{\prime}$ of $N$ is incident to two distinct transfer edges, with $v^{\prime}$ and $w^{\prime}$ as the other endpoints (where $u^{\prime},v^{\prime},w^{\prime}$ are subdivision nodes added when transforming $T$ to $N$). Since $N$ only adds transfers between minimal FA neighbors, this means that $v^{\prime}$ and $w^{\prime}$ are incomparable nodes of $\mathcal{T}_{N}$ (a special case arises when $v^{\prime}=w^{\prime}$, see the full proof for details). Roughly speaking, this means that two characters need to exchange material between the clades of $u^{\prime}$ and $v^{\prime}$, and also between the clades of $u^{\prime}$ and $w^{\prime}$. This is only possible in a galled-completion if the $u^{\prime}$ subtree contains at least two transfers, one exchanging with $v^{\prime}$ and the other with $w^{\prime}$. This creates two transfers below $u^{\prime}$ with external endpoints, contradicting the galled property by Lemma 2.2. ∎
@@
Suppose first that $v\neq w$. In that case, $C_{uv}$ and $C_{vw}$ are distinct as their FAs differ. By the second part of Lemma 3, the two characters $C_{uv}$ and $C_{uw}$ have a common FA $u$, and thus $v$ and $w$ are comparable. Suppose without loss of generality that $w\prec_{T}v$. We can further assume that $w$ is a minimal FA neighbor of $u$ (otherwise, we choose $w$ to be such a neighbor, which does not affect the presence of the edge between $u^{\prime}$ and $v^{\prime}$), which guarantees that $(w^{\prime},u^{\prime})\in E_{T}$. By the construction of $N$, the edge $(v^{\prime},u^{\prime})$ cannot be in $E_{T}$, as $u^{\prime}$ only receives transfers from its minimal neighbors. Therefore, the transfer between $u^{\prime}$ and $v^{\prime}$ is in the $(u^{\prime},v^{\prime})$ direction. As $v$ is not in a simple pair, this edge is present because $v$ has multiple FA neighbors and $u$ is minimal. This means that there is another character $C_{vz}$ with FAs $\{v,z\}$, where $z\neq u$. Using Lemma 3 in the same way as before, we get $u\prec_{T}z$. We now have two pairs of FAs $\{z,v\}$ and $\{u,w\}$, where $u\prec_{T}z$, $w$ is not comparable to $z$ (because $v$ is not), and $w\neq v$. By putting $\{z,v\}=\{a_{1},a_{2}\}$ and $\{u,w\}=\{b_{1},b_{2}\}$, we obtain a contradiction of the first part of Lemma 3.
So suppose that $v=w$, and hence $C_{uv}=C_{uw}$. This means that both edges $(u^{\prime},v^{\prime})$ and $(v^{\prime},u^{\prime})$ are present. This is not possible if $\{u,v\}$ is simple, so $u,v$ both have at least two FA neighbors, $v$ must be a minimal FA neighbor of $u$ and $u$ a minimal FA neighbor of $v$. Let $x\neq v$ be another FA neighbor of $u$. Then some character $C_{ux}$ has $\{u,x\}$ as FAs. Since $C_{uv}$ has FAs $\{u,v\}$, by the second part of Lemma 3, $x$ and $v$ must be comparable. Since $v$ is minimal, $v\prec_{T}x$. As $v$ also has other FA neighbors, by a symmetric argument, some character $C_{vy}$ has FAs $\{v,y\}$ with $u\prec_{T}y$. But the pair of FAs $\{y,v\},\{u,x\}$ has $u\prec_{T}y$, $x$ incomparable to $y$ ($x\prec_{T}y$ is not possible since $v$ would descend from $y$, and $y\prec_{T}x$ is not possible as $u$ would descend from $x$), and $v\neq x$, a contradiction of the first part of Lemma 3. ∎∎
@@
A tree $T$ on leafset $\mathcal{S}$ with character set $\mathcal{C}$ is galled-completable if and only if its redundancy-free network $N=(V,E_{S}\cup E_{T})$ is a galled tree.
@@
In the forward direction, if $T$ can be completed into a galled tree $N^{\prime}$ that explains $\mathcal{C}$, we can argue that for every transfer edge $(u,v)$ the is present in the redundancy-free network $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or in their support tree descendants. Roughly speaking, this means that the cycles of $N$ have corresponding cycles in $N^{\prime}$, and since $N^{\prime}$ is a galled tree, so is $N$. Conversely, if $N$ is galled, it explains $\mathcal{C}$ by construction. That is, for every character split into two clades in $T$, we either added an edge above the subtrees to connect them, or there is an edge that was added in the descendants due to some minimal FA neighbor, which allow meeting the connectivity requirements of Lemma 1. ∎
@@
($\Rightarrow$) Suppose that $T$ is galled-completable and let $N^{\prime}=(V^{\prime},E_{S}^{\prime}\cup E_{T}^{\prime})$ be a galled-completion for $T$ that explains $\mathcal{C}$. We need to show that $N$ is a galled tree. We assume that $N^{\prime}$ is obtained from $T$ by subdividing every edge once, then adding a set of transfer edges between subdivision nodes (note that if $N^{\prime}$ exists, such a galled-completion also exists). Since $N$ also subdivides every edge once, this allows us to assume that $N^{\prime}$ and $N$ have the same set of nodes, putting the nodes in correspondence without ambiguity.
@@
We first build an injective map, that associates each transfer edge $(u^{\prime},v^{\prime})$ of $N$ with a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$. Let $(u^{\prime},v^{\prime})\in E_{T}$. Note that by the construction of $N$, $u^{\prime},v^{\prime}$ are subdivision nodes of $\mathcal{T}_{N}$ and are parents of FAs $u,v$. Moreover, $u,v$ are FA neighbors, and thus there is some character $C$ with $\alpha_{T}(C)=\{u,v\}$. By Lemma 3, in $N^{\prime}$ there is a transfer edge $(x,y)$, where either $x\preceq_{\mathcal{T}_{N}}u^{\prime}$ and $y=v^{\prime}$, or $x\preceq_{\mathcal{T}_{N}}v^{\prime}$ and $y=u^{\prime}$. Either way, we map $(u^{\prime},v^{\prime})$ to $(x,y)$. We claim that no other transfer edge of $N$ can be mapped to $(x,y)$ in this manner. Suppose that some other transfer edge $(u^{\prime\prime},v^{\prime\prime})$ of $N$ maps to $(x,y)$. According to our map, $y=u^{\prime\prime}$ or $y=v^{\prime\prime}$, but either way, $y$ is incident to two transfer arcs in $N$, contradicting Lemma 3. Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$.
@@
Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its associated edge $(x,y)$ in $N^{\prime}$. Suppose that $(x,y)=(\tilde{u},v^{\prime})$ for some descendant $\tilde{u}$ of $u^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Consider the network $\tilde{N}$ obtained from $N^{\prime}$ by removing $(\tilde{u},v^{\prime})$ and inserting $(u^{\prime},v^{\prime})$ (which does nothing if $\tilde{u}=u^{\prime}$). Note that in $N^{\prime}$, the transfer edge $(\tilde{u},v^{\prime})$ belongs to a unique underlying cycle $H$ formed by the transfer edge, plus the paths from $\tilde{u}$ and from $v^{\prime}$ to the lowest common ancestor of $\tilde{u}$ and $v^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Since $u^{\prime}$ is on the path between that ancestor and $\tilde{u}$, in $\tilde{N}$ the incorporation of $(u^{\prime},v^{\prime})$ only creates an underlying cycle whose set of vertices is a subset of $H$. Since $H$ did not intersect with other underlying cycles in $N^{\prime}$, this new underlying cycle does not either, and it follows that $\tilde{N}$ is a galled tree as well. The same idea applies if $y=u^{\prime}$ instead.
@@
Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to replace every mapped transfer edge of $N^{\prime}$ by a transfer edge of $N$, one after another, while maintaining the property that no cycles intersect. Notice that because each $(u^{\prime},v^{\prime})\in E_{T}$ maps to a distinct transfer edge of $N^{\prime}$, every edge of $N$ is incorporated, and no edge of $N^{\prime}$ is “moved” twice by this process. Therefore, we obtain a galled tree whose set of transfer edges contains $E_{T}$. It then follows that $N$ is a galled tree as well.
@@
()($\Leftarrow$)( ⇐ ) Suppose that the redundancy-free network $N$ is a galled tree. We show that for every character in $C\in\mathcal{C}$ there exists a node in $N-F_{C}(N)$ that reaches every leaf in $C$ as required by Lemma 1. Note that if $|\alpha_{N}(C)|=1$, then $C$ forms a clade in $T$ and thus in $\mathcal{T}_{N}$. If $u$ is the FA node of $C$ in $T$, then $u$ is able to reach every element of $C$ in $N-F_{C}(N)$ since every support tree descendant of $u$ is also in $N-F_{C}(N)$, which is sufficient to explain $C$.
@@
Suppose that $|\alpha_{T}(C)|=2$. Let $u$ and $v$ be the FAs of $C$ in $T$, which are the roots of clades $C_{1}$ and $C_{2}$ such that $C=C_{1}\cup C_{2}$. Denote $p_{u}=p_{\mathcal{T}_{N}}(u)$ and $p_{v}=p_{\mathcal{T}_{N}}(v)$, where $p_{u}$ and $p_{v}$ are transfer nodes since every edge of $T$ was subdivided to produce $N$. Note that by definition, in $\mathcal{T}_{N}$ all the descendants of $p_{u}$ and of $p_{v}$ are in $C$, and thus both of them appear in $N-F_{C}(N)$. Suppose that one of the edges $(p_{u},p_{v})$ or $(p_{v},p_{u})$ is present. In the first case, $p_{u}$ is an origin for $C$ and in the second case, $p_{v}$ is an origin.
@@
So suppose that neither transfer edge is present. Then by construction, either $u$ or $v$ has at least two FA neighbors (otherwise, we would have added one of the two edges between their parent in an arbitrary direction and be in the previous case). Suppose without loss of generality that $u$ has at least two FA neighbors. Then $v$ is not one of its minimal FA neighbors, which means that there is some minimal FA neighbor $w$ of $u$ such that $w\prec_{T}v$. Let $p_{w}=p_{\mathcal{T}_{N}}(w)$ and note that $N$ contains the transfer edge $(p_{w},p_{u})$. Then, $v$ is an origin for $C$, since it is in $N-F_{C}(N)$, it can reach all the elements of $C$ that descend from itself in $\mathcal{T}_{N}$, and all those that descend from $u$ through the transfer edge $(p_{w},p_{u})$.
@@
3       Let $T^{\prime}$ be the tree that results from subdividing every edge in $T$, let $N=T^{\prime}$
4       // $FA(v)$ is a set that will keep track of FA neighbors present in $v$. for  $C\in\mathcal{C}$ do
@@ 25                         Mark $v$ as “possibly simple”
The previous lemma implies a polynomial time verification algorithm, detailed in Algorithm 1. We build the redundancy-free network and verify that it is a galled tree. We can calculate the set of FAs for each character and use this information to assign the FA neighbors to each node in $T$. Then, in a postorder traversal pick every node $v$ that has FA neighbors. Note that if those neighbors are not all comparable, then there exists no completion, since two outgoing transfers are necessary to explain them. When $v$ has multiple FA neighbors, we find the minimal one ($c_{min}$ in the algorithm) and add the corresponding transfer edge. If $c_{min}$ is the only FA neighbor of $v$, we “mark” $v$. If $c_{min}$ is also marked, then $\{v,c_{min}\}$ is a simple pair, and if not, then either $c_{min}$ will be marked later, or it has multiple FA neighbors and will create its own transfer.
@@
Let us remark that the complexity of the algorithm is dominated by the computation of the set of FA nodes. Assuming a traversal of $T$ for each $C\in\mathcal{C}$, this takes time $O(|V(T)||\mathcal{C}|)$. The rest of the algorithm only adds a time of $O(|V(T)|+|\mathcal{C}|)$. We leave the problem of computing FAs in linear time, if at all possible, for a future discussion.
@@
We begin by arguing that the algorithm is correct. The first for loop in Line 1 ensures that each character has at most two FAs as stated in Lemma 3. We claim that the second for loop will find all the required minimal transfers. For a fixed node $v$, we traverse its set of FA neighbors, noting that those are not necessarily comparable. However, if $v$ has two incomparable FA neighbors, then $v$ must also have at least two minimal incomparable FA neighbors. This then implies that the redundancy-free network $N$ will have two transfer edges incident to $p_{N}(v)$, a contradiction of Lemma 3. Thus, Line 1 allows us to rule out these cases. After we ensure that all the FA neighbors of $v$ are comparable we will find the minimum among them, which will allow us to add the desired transfer edge. For simple characters on the other hand, the verification in Line 1 allows us to add an arbitrary direction only when its unique neighbor is a marked node. The reason for this is that whenever the node is marked we know that there do not exist other FA neighbors to compare $c_{min}$ to.
Let us now argue the complexity. Let $T=(V,E)$ be the given tree. The complexity will be dominated by the first for loop. For a given character $C\in\mathcal{C}$, computing the FAs can be done in time $O(|V|)$. Thus the first for can be done in time $O(|\mathcal{C}||V|)$. On the other hand, the postorder traversal of $T$ that adds the transfer edges can be done in time $O(|V|+|\mathcal{C}|)$. This is because we traverse $O(|V|)$ nodes, and each $\{u,v\}$ relationship of FA neighborhood requires work at most twice throughout the algorithm, once for $u$ and once for $v$. Moreover, each character implies at most one FA relationship, the total work over the FA neighbor hood relationships is $O(|\mathcal{C}|)$. We note that checking for the incomparability or descendance of nodes can be achieved in constant time with standard pre-processing of the tree. Finally, the last verification on Line 1 can be done in time $O(|V|)$ (see e.g. [gusfield2014recombinatorics]). ∎∎

### v3

@@ 3 The Galled Completion Problem
Let $T$ be a tree on leafset $\mathcal{S}$ and character set $\mathcal{C}$. If $T$ is galled-completable, then in a redundancy-free network $N=(V_{T},E_{S}\cup E_{T})$ of $T$, every node is incident to at most one transfer edge.
@@
Suppose that some node $u^{\prime}$ of $N$ is incident to two distinct transfer edges, with $v^{\prime}$ and $w^{\prime}$ as the other endpoints. Here $u^{\prime},v^{\prime},w^{\prime}$ are subdivision nodes added when transforming $T$ to $N$, whose children in $\mathcal{T}_{N}$ are $u,v,w\in V(T)$, respectively. The presence of the two transfer edges implies that $\{u,v\}$ and $\{u,w\}$ are both sets of FAs of distinct characters. Moreover, since we only add edges between minimal FA neighbors, we get that $v$ and $w$ must be incomparable, as otherwise one of them would not be a minimal FA neighbor of $u$ (this requires case checking depending on the direction of the edges, along with a special case that arises when $v^{\prime}=w^{\prime}$, see the full proof for details). The pairs $\{u,v\}$ and $\{u,w\}$ of FAs then contradict Lemma 3, second part. ∎
@@
Suppose first that $v\neq w$. In that case, $C_{uv}$ and $C_{vw}$ are distinct as their FAs differ. By the second part of Lemma 3, the two characters $C_{uv}$ and $C_{uw}$ have a common FA $u$, and thus $v$ and $w$ are comparable. Suppose without loss of generality that $w\prec_{T}v$. We can further assume that $w$ is a minimal FA neighbor of $u$ (otherwise, we choose $w$ to be such a neighbor, which does not affect the presence of the edge between $u^{\prime}$ and $v^{\prime}$), which guarantees that $(w^{\prime},u^{\prime})\in E_{T}$. By the construction of $N$, the edge $(v^{\prime},u^{\prime})$ cannot be in $E_{T}$, as $u^{\prime}$ only receives transfers from the parents of the minimal FA neighbors of $u$. Therefore, the transfer between $u^{\prime}$ and $v^{\prime}$ is in the $(u^{\prime},v^{\prime})$ direction. As $v$ is not in a simple pair, this edge is present because $v$ has multiple FA neighbors and $u$ is minimal. This means that there is another character $C_{vz}$ with FAs $\{v,z\}$, where $z\neq u$. Using Lemma 3 in the same way as before, we get $u\prec_{T}z$ (and not $z\prec_{T}u$ because $u$ is a minimal FA neighbor of $v$). We now have two pairs of FAs $\{z,v\}$ and $\{u,w\}$, where $u\prec_{T}z$, $w$ is not comparable to $z$ (because $v$ is not), and $w\neq v$. By putting $\{z,v\}=\{a_{1},a_{2}\}$ and $\{u,w\}=\{b_{1},b_{2}\}$, we obtain a contradiction of the first part of Lemma 3.
So suppose that $v=w$. Recall that $u,v$ are the FAs of character $C_{uv}$, and by definition, this means that $C_{uv}=L(T(u))\cup L(T(v))$. Likewise, $C_{uw}=L(T(u))\cup L(T(w))$, and because $v=w$ we deduce that $C_{uv}=C_{uw}$ (also recall that we deal with sets of character, no so character is repeated). This means that both edges $(u^{\prime},v^{\prime})$ and $(v^{\prime},u^{\prime})$ are present. This is not possible if $\{u,v\}$ is simple, so $u,v$ both have at least two FA neighbors, $v$ must be a minimal FA neighbor of $u$ and $u$ a minimal FA neighbor of $v$. Let $x\neq v$ be another FA neighbor of $u$. Then some character $C_{ux}$ has $\{u,x\}$ as FAs. Since $C_{uv}$ has FAs $\{u,v\}$, by the second part of Lemma 3, $x$ and $v$ must be comparable. Since $v$ is minimal, $v\prec_{T}x$. As $v$ also has other FA neighbors, by a symmetric argument, some character $C_{vy}$ has FAs $\{v,y\}$ with $u\prec_{T}y$. But the pair of FAs $\{y,v\},\{u,x\}$ has $u\prec_{T}y$, $x$ incomparable to $y$ ($x\prec_{T}y$ is not possible since $v$ would descend from $y$, and $y\prec_{T}x$ is not possible as $u$ would descend from $x$), and $v\neq x$, a contradiction of the first part of Lemma 3. ∎∎
@@
A tree $T$ on leafset $\mathcal{S}$ with character set $\mathcal{C}$ is galled-completable if and only if any of its redundancy-free networks $N=(V,E_{S}\cup E_{T})$ is a galled tree.
@@
In the forward direction, if $T$ can be completed into a galled LGT network $N^{\prime}$ that explains $\mathcal{C}$, we can argue that for every transfer edge $(u,v)$ that is present in a redundancy-free network $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or $u$ is a support tree descendant of the sending end of that transfer edge in $N$. Roughly speaking, this means that the cycles of $N$ have corresponding cycles in $N^{\prime}$, and since $N^{\prime}$ is a galled tree, so is $N$. Conversely, if $N$ is galled, it explains $\mathcal{C}$ by construction. That is, for every character split into two clades in $T$, we either added an edge above the subtrees to connect them, or there is an edge that was added in the descendants due to some minimal FA neighbor, which allow meeting the connectivity requirements of Lemma 1. ∎
@@
($\Rightarrow$) Suppose that $T$ is galled-completable and let $N^{\prime}=(V^{\prime},E_{S}^{\prime}\cup E_{T}^{\prime})$ be a galled-completion for $T$ that explains $\mathcal{C}$. Let $N=(V,E_{S}\cup E_{T})$ be any redundancy-free network of $T$. We need to show that $N$ is a galled tree. We assume that $N^{\prime}$ is obtained from $T$ by subdividing every edge once, then adding a set of transfer edges between subdivision nodes (note that if $N^{\prime}$ exists, such a galled-completion also exists). Since $N$ also subdivides every edge once, this allows us to assume that $N^{\prime}$ and $N$ have the same set of nodes, putting the nodes in correspondence without ambiguity.
@@
We first build an injective map, that associates each transfer edge $(u^{\prime},v^{\prime})$ of $N$ with a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$, and such that $x$ descends from the node of $\{u^{\prime},v^{\prime}\}\setminus\{y\}$ in $\mathcal{T}_{N^{\prime}}$. Let $(u^{\prime},v^{\prime})\in E_{T}$. Note that by the construction of $N$, $u^{\prime},v^{\prime}$ are subdivision nodes of $\mathcal{T}_{N}$ and, in that support tree, they are parents of nodes $u,v\in V(T)$. Moreover, $u,v$ are FA neighbors, and thus there is some character $C$ with $\alpha_{T}(C)=\{u,v\}$. By Lemma 3, in $N^{\prime}$ there is a transfer edge $(x,y)\in E^{\prime}_{T}$, where either $x\preceq_{\mathcal{T}_{N}}u^{\prime}$ and $y=v^{\prime}$, or $x\preceq_{\mathcal{T}_{N}}v^{\prime}$ and $y=u^{\prime}$. Either way, we map $(u^{\prime},v^{\prime})$ to $(x,y)$ (note that our desired properties on the map hold). We claim that no other transfer edge of $N$ can be mapped to $(x,y)$ in this manner. Suppose that some other transfer edge $(u^{\prime\prime},v^{\prime\prime})$ of $N$ maps to $(x,y)$. According to our map, $y=u^{\prime\prime}$ or $y=v^{\prime\prime}$, but either way, $y$ is incident to two transfer edge $(u^{\prime},v^{\prime})$ and $(u^{\prime\prime},v^{\prime\prime})$ in $N$, contradicting Lemma 3. Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$ and the mapping is indeed injective.
@@
Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its associated edge $(x,y)\in E^{\prime}_{T}$ in $N^{\prime}$. According the the properties of our map, suppose first that $(x,y)=(\tilde{u},v^{\prime})$ for some descendant $\tilde{u}$ of $u^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Consider the network $\tilde{N}$ obtained from $N^{\prime}$ by removing $(\tilde{u},v^{\prime})$ and inserting $(u^{\prime},v^{\prime})$ (which does nothing if $\tilde{u}=u^{\prime}$). Note that in $N^{\prime}$, the transfer edge $(\tilde{u},v^{\prime})$ belongs to a unique underlying cycle $H$ formed by the transfer edge, plus the paths from $\tilde{u}$ and from $v^{\prime}$ to the lowest common ancestor of $\tilde{u}$ and $v^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Since $u^{\prime}$ is on the path between that ancestor and $\tilde{u}$, in $\tilde{N}$ the incorporation of $(u^{\prime},v^{\prime})$ only creates an underlying cycle whose set of vertices is a subset of $H$. Since $H$ did not intersect with other underlying cycles in $N^{\prime}$, this new underlying cycle does not either, and it follows that $\tilde{N}$ is a galled tree as well. The same idea applies if $y=u^{\prime}$ and $x\preceq_{\mathcal{T}_{N^{\prime}}}v^{\prime}$ instead, that is, we can remove $(x,y)$ from $N^{\prime}$ and add $(u^{\prime},v^{\prime})$ and the result $\tilde{N}$ is still galled.
@@
Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to insert into $N^{\prime}$ every transfer edge of $N$, one after another, while maintaining the property that no cycles intersect. Notice that because each $(u^{\prime},v^{\prime})\in E_{T}$ maps to a distinct transfer edge of $N^{\prime}$, every edge of $N$ is incorporated, and no edge of $N^{\prime}$ is “moved” twice by this process. Therefore, we obtain a galled tree whose set of transfer edges contains $E_{T}$. It then follows that $N$ is a galled tree as well.
@@
()($\Leftarrow$)( ⇐ ) Take some redundancy-free network $N$ of $T$ and assume that $N$ is a galled tree. We show that for every character in $C\in\mathcal{C}$ there exists a node in $N-F_{C}(N)$ that reaches every leaf in $C$ as required by Lemma 1. Note that if $|\alpha_{N}(C)|=1$, then $C$ forms a clade in $T$ and thus in $\mathcal{T}_{N}$. If $u$ is the FA node of $C$ in $T$, then $u$ is able to reach every element of $C$ in $N-F_{C}(N)$ since every support tree descendant of $u$ is also in $N-F_{C}(N)$, which is sufficient to explain $C$.
@@
Suppose that $|\alpha_{T}(C)|=2$. Let $u$ and $v$ be the FAs of $C$ in $T$, which are the roots of clades $C_{1}$ and $C_{2}$ such that $C=C_{1}\cup C_{2}$. Denote $p_{u}=p_{\mathcal{T}_{N}}(u)$ and $p_{v}=p_{\mathcal{T}_{N}}(v)$, where $p_{u}$ and $p_{v}$ are transfer nodes since every edge of $T$ was subdivided to produce $N$. Note that by definition, in $\mathcal{T}_{N}$ all the descendants of $p_{u}$ and of $p_{v}$ are in $C$, and thus both $p_{u}$ and $p_{v}$ appear in $N-F_{C}(N)$. Suppose that one of the transfer edges $(p_{u},p_{v})$ or $(p_{v},p_{u})$ is present in $N$. In the first case, $p_{u}$ is an origin for $C$ and in the second case, $p_{v}$ is an origin.
@@
So suppose that neither transfer edge is present. Because $u$ and $v$ are FA neighbors, this is only possible if either $u$ or $v$ has at least two FA neighbors (otherwise, we would have added one of the two edges between their parent in an arbitrary direction and be in the previous case). Suppose without loss of generality that $u$ has at least two FA neighbors. Then $v$ is not one of its minimal FA neighbors, which means that there is some minimal FA neighbor $w$ of $u$ such that $w\prec_{T}v$. Let $p_{w}=p_{\mathcal{T}_{N}}(w)$ and note that $N$ contains the transfer edge $(p_{w},p_{u})$. Then, $v$ is an origin for $C$, since it is in $N-F_{C}(N)$, it can reach all the elements of $C$ that descend from itself in $\mathcal{T}_{N}$, and all those that descend from $u$ through the transfer edge $(p_{w},p_{u})$.
@@
3       Let $T^{\prime}$ be obtained from $T$ by subdividing every edge, let $N=T^{\prime}$
4       Initialize $FA(v)=\emptyset$ for all $v\in V(N)$, used to store the FA neighbors of $v$.
5       for  $C\in\mathcal{C}$ do
@@ 26                         Mark $v$ as “possibly simple”
The previous lemma implies a polynomial time verification algorithm, detailed in Algorithm 1. We build a redundancy-free network and verify that it is a galled tree. We can calculate the set of FAs for each character and use this information to assign the FA neighbors to each node in $T$. Then, in a postorder traversal pick every node $v$ that has FA neighbors. Note that if those neighbors are not all comparable, then there exists no completion, since two outgoing transfers are necessary to explain them. When $v$ has multiple FA neighbors, we find the minimal one ($c_{min}$ in the algorithm) and add the corresponding transfer edge. If $c_{min}$ is the only FA neighbor of $v$, we “mark” $v$. If $c_{min}$ is also marked, then $\{v,c_{min}\}$ is a simple pair, and if not, then either $c_{min}$ will be marked later, or it has multiple FA neighbors and will create its own transfer.
@@
Let us remark that the complexity of the algorithm is dominated by the computation of the set of FA nodes. Assuming a traversal of $T$ for each $C\in\mathcal{C}$, this takes time $O(|V(T)||\mathcal{C}|)$. The rest of the algorithm only adds a time of $O(|V(T)|+|\mathcal{C}|)$. Note that verifying whether a given network is a galled tree, as required in Line 1 can be done in time $O(|V|)$. Indeed, in a galled tree $N=(V,E)$ the number of edges is $O(|V|)$. Moreover, galled trees are the (binary) networks whose biconnected components contain at most one reticulation node (see e.g. [gusfield2014recombinatorics, Chapter 8], a biconnected component is a subgraph that cannot be disconnected by removing a single vertex, and a reticulation is a node of indegree $2$). We can thus use Tarjan’s algorithm to find biconnected components in linear time [tarjan1972depth], then verify that each of them contains at most one reticulation node. Since those components are vertex-disjoint, going through all of them takes time $O(|V|)$. We leave the problem of computing FAs in linear time, if at all possible, for a future discussion.
@@
We begin by arguing that the algorithm is correct. The first for loop in Line 1 ensures that each character has at most two FAs as stated in Lemma 3. We claim that the second for loop will find all the required minimal transfers. For a fixed node $v$, we traverse its set of FA neighbors, noting that those are not necessarily comparable. However, if $v$ has two incomparable FA neighbors, then $v$ must also have at least two minimal incomparable FA neighbors. This then implies that a redundancy-free network $N$ will have two transfer edges incident to $p_{N}(v)$, a contradiction of Lemma 3. Thus, Line 1 allows us to rule out these cases. After we ensure that all the FA neighbors of $v$ are comparable we will find the minimum among them, which will allow us to add the desired transfer edge. For simple characters on the other hand, the verification in Line 1 allows us to add an arbitrary direction only when its unique neighbor is a marked node. The reason for this is that whenever the node is marked we know that there do not exist other FA neighbors to compare $c_{min}$ to.
Let us now argue the complexity. Let $T=(V,E)$ be the given tree. The complexity will be dominated by the first for loop. For a given character $C\in\mathcal{C}$, computing the FAs can be done in time $O(|V|)$. Thus the first for can be done in time $O(|\mathcal{C}||V|)$. On the other hand, the postorder traversal of $T$ that adds the transfer edges can be done in time $O(|V|+|\mathcal{C}|)$. This is because we traverse $O(|V|)$ nodes, and each $\{u,v\}$ relationship of FA neighborhood requires work at most twice throughout the algorithm, once for $u$ and once for $v$. Moreover, each character implies at most one FA relationship, the total work over the FA neighbor hood relationships is $O(|\mathcal{C}|)$. We note that checking for the incomparability or descendance of nodes can be achieved in constant time with standard pre-processing of the tree. Finally, the last verification on Line 1 can be done in time $O(|V|)$ (see e.g. [gusfield2014recombinatorics] and main text). ∎∎